OpenStudy (anonymous):

A subway train starts from rest at a station and accelerates at a rate of 1.60 m/s2 for 14.0s. It runs constant speed for 70.0s and slows down at a rate of 3.50 m/s2 until it stops at the next station find the total distance covered.

OpenStudy (anonymous):

constant speed = acceleration times time = 1.6*14 = 22.4 m/s area of first triangle = 1/2*(1.6)* 14^2 = 156.8 m area of rectangle = 1.6 * 14 *70 = 156.8 m time in the last phase : t = 1.6*14/3.5 s area of the last triangle = 1/2 *1.6 * 14 * 1.6 * 14 / 3.5 = 71.68 m Total distance = 156.8 + 156.8 + 71.68 = 385.28 m